OpenStudy (timinthor):

A line passes through (2, −1)(2, −1) and (4, 5)(4, 5). What is the equation of the line?

OpenStudy (studygurl14):

First find the slope by using the slope formula a line goes through two points \((x_1,y_1)\) and \((x_2,y_2)\) Slope=\(\Large\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (timinthor):

I did that. I got 3 for the slope.

OpenStudy (mrnood):

why have you written the points twice?

OpenStudy (studygurl14):

@MrNood He hasn't. Openstudy sometimes copies it for some weird reason

OpenStudy (studygurl14):

Good job, it is 3 @timinthor

OpenStudy (studygurl14):

Now, plug that into the equation y=mx + b m = slope Use a given point to plug in for y and x and solve for b

OpenStudy (timinthor):

So y=3x+b but how do i find b? @StudyGurl14

OpenStudy (studygurl14):

Use a given point. (2,-1) for example Remember that a point is (x,y) Plug in for x and y in the equation and solve for b

OpenStudy (timinthor):

So -1=3x+2?

OpenStudy (timinthor):

Do you solve like a normal equation? @StudyGurl14 If so, -1=3x+2 -2 -2 -3=3x -3/3 x=-1? But it asks for the equation. None of the answers look like that though. The answer choices are: A. −3x+5y=−13 B. −3x+y=−7 C. −3x+5y=13 D. −3x+y=17

OpenStudy (studygurl14):

No, you plugged in 2 for b, not x. y = mx + b y = 3x + b (-1) = 3(2) + b Now solve for b

OpenStudy (timinthor):

-7? @StudyGurl14

OpenStudy (studygurl14):

yep. so you get y = 3x - 7

OpenStudy (timinthor):

Thanks!

OpenStudy (studygurl14):

:)